how to find directrix of 9x2+12xy+4y2-98x-74y+300=0 is

Grade:12th pass

2 Answers

Aditya Gupta
4 years ago
so this is an easy ques and i am getting equation of directrix as 2x – 3y+5=0
reason is coz the above equation can be transformed into:
(x – 3)^2 + (y – 4)^2 = (2x – 3y+5)^2/(2^2+(-3)^2)
this equation is the “focus-directrix” form of the parabola. 
hence we can directly conclude that directrix would be 2x – 3y+5=0
Sai Anil lahane
one year ago
the equation 9x ^ 2 + 12xy + 4y ^ 2 - 98x - 74y + 300 = 0 represents (a) a pair of parallel lines (d) a hyperbola